Circular No. 5603

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

RX J2117.1+3412 T. K. Watson, Department of Astronomy, University of Texas, Austin,

reports: “We have discovered that the white dwarf RX J2117.1+3412 (α = 21^{h}17^{m}07^{s}, δ = +34^{o}12′.4, equinox 2000.0; B = 13) is a pulsating variable star. K. Werner (1992, 8th European Workshop on White Dwarf Stars, Leicester) reports that the object is a ROSAT source and a member of the PG 1159–035 spectroscopic class, so this is the fifth known DOV star and by far the brightest member of its class. We observed the star with a high-speed, two-star photometer on several nights, using Hamamatsu R647-04 (bialkali) phototubes in unfiltered light on the 2.1-m telescope at the McDonald Observatory. We find pulsations with a peak-to-peak amplitude of about 0.03 mag and a quasiperiod of 850 s. Each of three 90-120-min runs during Aug. 20-28 has a power spectrum that shows significant bands of power between 500 and 900 s, the amplitude of the largest peak in each power spectrum being 0.007 mag. There are indications that the bands of power observed are made up of closely spaced periodicities, similar to those seen in other DOV stars, but our runs are not yet long enough to resolve them. Observers interested in studying this star should send an e-mail request for a finding chart to tkw@bullwinkle.as.utexas.edu and include their fax number.”

1992 LR

Continuation to the ephemeris on IAUC 5581:

1992TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  V
Sept. 5 22^{h}16.^{m}07 + 4^{o}11.^{′}7 0.162 1.165 165.^{o}1 12.^{o}9 15.1
10 22 27.64 + 3 25.8 0.183 1.185 165.6 12.2 15.4
15 22 37.36 + 2 42.7 0.206 1.206 165.1 12.4 15.7
20 22 45.72 + 2 04.0 0.232 1.228 163.5 13.4 16.1
25 22 53.12 + 1 30.7 0.260 1.252 161.2 15.0 16.4
30 22 59.93 + 1 03.7 0.291 1.277 158.3 16.8 16.7
Oct. 5 23 06.39 + 0 43.4 0.325 1.302 155.2 18.8 17.1
10 23 12.67 + 0 30.0 0.362 1.328 151.8 20.8 17.4
15 23 18.88 + 0 23.1 0.401 1.355 148.4 22.7 17.8
20 23 25.09 + 0 22.3 0.443 1.383 145.0 24.4 18.1
25 23 31.37 + 0 27.3 0.489 1.410 141.6 26.0 18.4

1992 September 3 Brian G. Marsden